You find an ancient mammal frozen in a glacier. Yo Please help and explain
Answer #1
  • GC content % is calculated as
  • G + C/ A+T+G+C = 66 = A+17+66
  • A = 17 +66-66 = 17
  • A+T/G+C is the ratio of AT and GC
  • Length of the sequence is directly proportional to the G+C content.
  • The amounts of A and T are always equal.
  • Thus if T = 17%, then A is also 17%
  • Thus G is 33% and C is 33%
  • So, here frozen specimen sloths have 34% of AT and 66% of GC. This shows that the buried frozen mammal is a member of modern sloths but not an extinct species.